# LASeR: Learning to Adaptively Select Reward Models with Multi-Armed Bandits

### v1

@@ Abstract
Reward Models (RMs) play a crucial role in aligning large language models (LLMs) with human preferences, enhancing their performance by ranking outputs during inference or iterative training. However, the degree to which an RM generalizes to new tasks is often not known a priori. For instance, some RMs may excel at scoring creative writing, while others specialize in evaluating math reasoning. Therefore, using only one fixed RM while training LLMs can be suboptimal. Moreover, optimizing LLMs with multiple RMs simultaneously can be prohibitively computationally-intensive and challenging due to conflicting signals from different RMs, potentially degrading performance. To address these challenges, we introduce LASeR (Learning to Adaptively Select Rewards), which iteratively trains LLMs using multiple RMs, selecting and utilizing the most well-suited RM for each instance to rank outputs and generate preference data, framed as a multi-armed bandit problem. Our empirical results on commonsense and math reasoning tasks demonstrate that LASeR can boost iterative LLM optimization by optimizing for multiple RMs, improving the absolute average accuracy of Llama-3-8B over three datasets by $2.67\%$ over training with ensemble RM scores while also showing superior training efficiency (e.g., a 2×2$\times$2 × speedup). Moreover, on WildChat, a benchmark of instruction-following prompts in open-form generation, we find that using Llama-3-8B LASeR leads to a $71.45\%$ AlpacaEval win rate over sequentially optimizing multiple RMs. Extending to long-context generation tasks, we find that on Llama-3-8B, LASeR achieves an average improvement of $2.64$ F1 points on single-document QA tasks and $2.42$ F1 points on multi-document QA over random RM selection when used with best-of-$n$ sampling. Our analysis shows that LASeR is robust to noisy rewards and generalizes to multiple settings. Finally, we demonstrate that LASeR’s RM selection changes depending on the underlying task or instance, and we verify the presence of conflicting preferences from multiple RMs, which can be mitigated using LASeR. [^1] Code: https://github.com/duykhuongnguyen/LASeR-MAB
**footnotetext: Equal Contribution
@@ 1 Introduction
When comparing two responses, human preferences often differ depending on factors like the underlying task, who the annotators are (Santurkar et al., 2023; Ahmadian et al., 2024), and how preferences are elicited (Bansal et al., 2024). Therefore, models of preference data are also likely to differ, and might include noise as well as any biases contained in the preference data used to train them. This can pose a problem when using such models as “reward models” (RMs) to align large language models (LLMs) to human preferences using reinforcement learning with human feedback (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022). Recent work has focused on aligning LLMs through iterative training, using reward models as proxies for human judgment (Gulcehre et al., 2023), leveraging the LLM to act as an implicit RM or judge (Yuan et al., 2024; Chen et al., 2024b), or using the gold answer to compute a reward (Pang et al., 2024). Under this paradigm, there are three stages to training LLMs: (i) generating multiple responses to a query from an LLM; (ii) scoring the responses with an RM to create preference data with better and worse responses; and (iii) using model-generated preference data to further train the LLM. Note that for most domains, the gold reward is not readily available, making the quality of the RM or the degree to which it reflects human preferences (i.e., the gold reward) crucial to improving LLM performance. Indeed, several prior efforts aim to train new RMs that better reflect human preferences (Lambert et al., 2024).
@@
However, selecting one reward model to guide LLM training can be suboptimal for three main reasons: (1) A single RM may not be generalized to heterogeneous sets of examples. RMs are typically designed to reflect specific objectives and may be trained on offline preference datasets. Thus, an RM that performs well on one dataset or domain may not generalize effectively to others, leading to misaligned outputs across different tasks or domains (Kirk et al., 2023; Chen et al., 2024a; Casper et al., 2023; Gao et al., 2023). For instance, creativity plays a key role in evaluating the quality of a story, whereas correctness is more important in scoring math solutions. (2) RM performance leaderboards (e.g., Lambert et al. (2024)) that rely on human-annotated preferences can have unreliable rankings due to the presence of incorrect and ambiguous preferences (Yu et al., 2024; Hejna et al., 2023). (3) Lastly, over-optimization on a particular RM can lead to reward hacking (Skalse et al., 2022; Rafailov et al., 2024a), resulting in minimal gain or even drops in downstream performance.
@@
To mitigate these issues, a prevalent approach is to ensemble multiple reward models (Coste et al., 2023; Eisenstein et al., 2023; Zhang et al., 2024; Ramé et al., 2024). However, these methods also come with significant challenges: as RMs are typically based on LLMs, training with multiple RMs often requires loading and managing several large models simultaneously, which can be computationally expensive, becoming infeasible as models increase in size. Moreover, aggregating multiple RM scores together is susceptible to noisy rewards or conflicting preferences from RMs, especially RMs that are not well-suited for the specific task (Rita et al., 2024). This in turn, can degrade the quality of the preference data, leading to low-quality updates during training (Wang et al., 2024a). Finally, manually selecting a subset of RMs to combine is a labor-intensive process that involves training many different variants on a combinatorially large set of RM groupings. This underscores the need for more efficient methods that efficiently and robustly optimize LLMs using multiple RMs.
@@
In this work, we introduce Learning to Adaptively Select Rewards (LASeR), that, given a set of RMs, adaptively and efficiently selects a suitable RM for each instance by casting selection as a multi-armed bandit problem (Vermorel & Mohri, 2005; Audibert et al., 2009). Specifically, during training, the RM (arm) is chosen dynamically based on contextual information about the model’s performance and past interactions. The LLM is then fine-tuned based on the RM-annotated data, and the bandit’s parameters are updated accordingly to reflect the performance of the LLM after training on preference data annotated using selected RM (see Fig. 1). By design, LASeR’s adaptive instance-level RM selection (c.f. Sec. 3) addresses the three shortcomings of choosing one reward model (lack of generalization, unreliable rankings, and over-optimization) and outperforms using the same RM across all instances, yielding higher downstream performance and better generalization. Moreover, unlike previous multi-RM methods that require simultaneously loading and running multiple RMs (Ramé et al., 2024; Coste et al., 2023), our method selects one reward model at a time (Sec. 4). This makes the training more efficient and improves performance by allowing the model to adaptively focus on the most suitable RM for each specific instance or phase of training.
@@
Empirically, we demonstrate the effectiveness of LASeR for iteratively training LLMs using multiple RMs on three broad domains: reasoning, instruction-following in text generation, and long-context understanding (Sec. 4.2). We show that on reasoning benchmarks such as StrategyQA (Geva et al., 2021) (testing commonsense reasoning) and GSM8K (Cobbe et al., 2021) (testing math reasoning), LASeR with Llama-3-8B outperforms training with the single best RM by $1.45\%$, and jointly optimizing with ensemble (average) RM scores by $2.67\%$ in absolute accuracy averaged over datasets. LASeR is also effective on general instruction-following: we show that using LASeR with four strong 7B RMs from RewardBench to finetune Llama-3-8B on a subset of WildChat (Zhao et al., 2024) beats LLMs trained with the best RM in the ensemble and a sequential baseline, with $56.34\%$ and $71.45\%$ win rates (respectively) on length-controlled AlpacaEval (Dubois et al., 2024). Moreover, our results show the effectiveness of LASeR’s RM selection strategy at inference-time on long-form generation tasks; on LongBench (Bai et al., 2022), we find LASeR beats random RM selection baseline by $2.64$ F1 points on single-document QA tasks and $2.42$ F1 points on multi-document QA when using best-of-$n$ sampling for Llama-3-8B. Our analysis reveals that LASeR is more efficient than sequential multi-RM and RM ensemble baselines in terms of training time (wall-clock hours) by a factor of 3×3$\times$3 ×, and 2×2$\times$2 ×, respectively while being more robust to noisy rewards and conflicting preferences from multiple RMs (Sec. 5). Finally, we demonstrate that LASeR effectively selects RMs based on the underlying instance and generalizes to multiple settings, including out-of-distribution datasets, different training loss functions, etc.
@@ 2 Related work
Multiple Reward Ensembles.
Training large language models (LLMs) with multiple reward functions is an emerging research area focused on aligning model outputs with complex objectives that require diverse evaluation metrics. One prior approach involves using ensembles of multiple rewards (Ramé et al., 2024; Wu et al., 2024; Coste et al., 2023; Zhang et al., 2024; Wang et al., 2024b; Jang et al., 2023; Eisenstein et al., 2023). Unlike these methods, which train multiple RMs or use multiple RMs during LLM training, LASeR selects only a single pretrained RM at each LLM training step. Not only is optimizing for one RM at a time more efficient, but it also avoids the problem of having conflicting rewards (Rita et al., 2024). In Sec. 4.2, we show that LASeR outperforms an offline variant of optimizing LLMs with an ensemble of RMs (consistent with the mean optimization setting from Coste et al. (2023)). Another line of work uses mixture-of-experts (MoE) techniques for training an RM from multiple interpretable objectives (Wang et al., 2024c) or jointly training task-specific RMs and a sparse MoE router (Quan, 2024). Instead of relying on static datasets with human-annotated attributes for RM training (as Wang et al. (2024c) does), we employ existing RMs to train LLMs on its own generations, a strategy that has been shown to be more effective (Ivison et al., 2024). Unlike Quan (2024), who jointly train multiple RMs and sparse MoE router (requiring loading all models at once), LASeR implicitly and efficiently learns a sparse router and does not need to train RMs. LASeR performs RM selection via multi-armed bandits over existing off-the-shelf RMs with strong performance on leaderboards (Lambert et al., 2024).
Multi-Armed Bandits (MABs).
There is a long-standing history of using multi-armed bandit algorithms for diverse applications in machine learning spanning online advertising (Chen et al., 2013), recommendation systems (Li et al., 2010), hyperparameter optimization (Li et al., 2018), curriculum learning (Graves et al., 2017), with some recent work at the intersection of MABs and language models (Pasunuru et al., 2020; Krishnamurthy et al., 2024; Dwaracherla et al., 2024, inter alia). In the realm of RLHF specifically, Dwaracherla et al. (2024) use double Thomson Sampling (Wu & Liu, 2016) to select which of the sampled responses should be annotated and paired using a single fixed RM, improving LLM performance. In contrast, LASeR first selects which RM (i.e., model of preferences) should be used to annotate LLM’s responses to a query and then creates multiple preference pairs from these responses. Pasunuru et al. (2020) optimize text generation models for different evaluation metrics such as ROUGE-L and BLEU via policy gradients (Williams, 1992) over existing (static) question and data-to-text generation datasets. In contrast, LASeR adopts an iterative training recipe and improves downstream generation performance across a wide range of tasks from instruction-following to math and commonsense reasoning by selecting relevant RMs and scoring the LLM’s own generated responses without access to true rewards, i.e., gold labels, and without optimizing for the downstream evaluation metric.
Iterative LLM Training.
Recent works on training LLMs incorporate reinforcement learning with human feedback (RLHF) due to its effectiveness in improving instruction following the ability of LLMs over their pretrained counterparts (Ouyang et al., 2022; Bai et al., 2022; Touvron et al., 2023). The standard RLHF training pipeline, which comprises finetuning the LLM on static datasets with human-annotated preferences, is bottlenecked by the size and quality of annotated preference data as well as the limited effectiveness of off-policy optimization (Xu et al., 2023; Xiong et al., 2024; Yuan et al., 2024; Guo et al., 2024). To remedy this, recent work focuses on training LLMs iteratively, scoring the LLM’s own generations to create feedback data for RLHF. This line of work obtains scores either from gold labels (Singh et al., 2023; Pang et al., 2024), from a single RM (Gulcehre et al., 2023), or from the generating LLM (Yuan et al., 2024; Chen et al., 2024b). In contrast, we take advantage of the abundance of publicly-available RMs and growing interest in developing RMs (Lambert et al., 2024) by using multiple RMs. This has several advantages over past work: it works in cases where gold labels are not available (e.g., generation tasks), deals with the three issues associated with using a single RM (unreliable rankings, lack of generalizability, and over-optimization), reduces the burden on the user to pick the right RM, and avoids problems stemming from the inability of LLMs to judge their own responses for certain domains (Huang et al., 2023).
@@ 3 LASeR: Learning to Adaptively Select Rewards
In this section, we describe LASeR in detail. First, we expand on the training pipeline (in one iteration) with a general reward function (Sec. 3.1). Then, in Sec. 3.2, we describe how LASeR dynamically selects an RM from a set of multiple RMs using MAB algorithms, i.e., how we dynamically assign the reward function for a given instance or batch of instances. Finally, in Sec. 3.3, we describe the overall training setup across iterations and specifically how we update the parameters of the MAB at the end of each iteration. A detailed illustration of LASeR is shown in Fig. 1.
Figure 1: Overview of LASeR. Given the query, the multi-armed bandit selects an RM depending on the underlying query and the bandit’s parameters (based on the usage of each RM and the expected MAB reward). At iteration $m$, the LLM generates multiple responses that are scored based on the selected RM for that query, scoring each response. These responses are ranked into preference pairs, which are then used to fine-tune the model. The same train loss $\mathcal{L}^{m}$ is used to update the parameters of the LLM as well as the MAB for the next iteration, making the entire pipeline iterative.
@@ 3.1 Training LLMs using a Reward Function
Following Yuan et al. (2024) and Pang et al. (2024), we adopt an iterative training pipeline to finetune the LLM for $M$ iterations. Let $\pi_{m}$ be the LLM at iteration $m$; we assume that we start from an initial pretrained model $\pi_{0}$. Let $\mathcal{D}=\{x_{1},x_{2},\dots,x_{N}\}$ represent the training inputs, where $x_{i}$ is an input query or prompt. Corresponding to each input query $x_{i}$, we sample a set of $n$ responses from the LLM at the current $m^{\text{th}}$ iteration as $\mathbf{y}_{i}=\{y^{1}_{i},y^{2}_{i},\dots,y^{n}_{i}\}\sim\pi_{m}(y|x_{i})$. Let $R^{\star}\!:\!(y_{i}^{j}|x_{i})\rightarrow\mathbb{R}$ be a reward function that can score an LLM generated response $y_{i}^{j}$ to a query $x_{i}$ based on how well it aligns with specific task objectives or instructions. Note that $R^{\star}(.)$ can be any reward function and may correspond to a single RM, one of the multiple RMs selected by the MAB (as in our case), or even the true reward.
Generating Preference Pairs.
@@
We evaluate each response $y_{i}^{j}$ using the reward function $R^{\star}(y_{i}^{j}|x_{i})$. By comparing the rewards assigned to different responses, we can form preference pairs $(y_{i}^{w},y_{i}^{l})$, where $y_{i}^{w}$ is preferred over $y_{i}^{l}$ if $R^{\star}(y_{i}^{w}|x_{i})>R^{\star}(y_{i}^{l}|x_{i})$, thereby building a preference dataset: [^2] Following Pang et al. (2024), in our experiments, we randomly sample $P\!=\!10$ pairs corresponding to each prompt $x_{i}$. For brevity, we omit this in the notation of $\mathcal{D}_{\mathrm{pref}}$; but in our setting $|\mathcal{D}_{\mathrm{pref}}|=P\times|\mathcal{D}|$.
$\mathcal{D}_{\mathrm{pref}}=\{(x_{i},y_{i}^{w},y_{i}^{l})\mid x_{i}\in\mathcal{D},R^{\star}(y_{i}^{w})>R^{\star}(y_{i}^{l})\}.$
Training Loss Function ($\mathcal{L}^{m}$).
@@
In each iteration, we fine-tune the model using the generated preference dataset $\mathcal{D}_{\mathrm{pref}}$, resulting in $M$ models $\pi_{1},\pi_{2},\dots,\pi_{M}$. Specifically, we update the model using the DPO loss (Rafailov et al., 2024b) for learning from the preference pairs. In this work, we use the following loss functions for training the LLM at iteration $m$:
@@
$\begin{split}\mathcal{L}^{m}_{\textsc{dpo}}(\pi_{m})\!&=\!-\mathbb{E}_{(x_{i},y_{i}^{w},y_{i}^{l})\sim\mathcal{D}_{\mathrm{pref}}}\left[\log\sigma\left(\beta\log\frac{\pi_{m}(y_{i}^{w}\!\mid\!x)}{\pi_{m-1}(y_{i}^{w}\mid x)}-\beta\log\frac{\pi_{m}(y_{i}^{l}\!\mid\!x)}{\pi_{\scriptstyle m-1}(y_{i}^{l}\!\mid x)}\right)\right]\\ \mathcal{L}^{m}_{\textsc{nll}}(\pi_{m})\!&=\!-\mathbb{E}_{(x_{i},y_{i}^{w})\sim\mathcal{D}_{\mathrm{pref}}}\left[\frac{\log\pi_{m}(y_{i}^{w}\!\mid\!x_{i})}{|y_{i}^{w}|}\right],\end{split}$ (1)
@@
where $\pi_{m}$, and $\pi_{m-1}$ denotes the LLM in the current iteration $m$ and the previous iteration $m-1$ (used as the reference model in DPO loss). Following Yuan et al. (2024), we use the standard DPO loss for instruction-finetuning. Following Pang et al. (2024), we use the NLL loss on the preferred responses as an additional regularizer for reasoning tasks, i.e., $\mathcal{L}^{m}\!=\!\mathcal{L}^{m}_{\textsc{dpo}}+\mathcal{L}^{m}_{\textsc{nll}}$. In Appendix C, we show that LASeR outperforms baselines irrespective of the choice of the loss function $\mathcal{L}^{m}$.

### v2

@@ Abstract
Reward Models (RMs) are crucial to aligning large language models (LLMs), but the degree to which an RM specialized to one task (e.g. writing) generalizes to new tasks (e.g. math) is often not known a priori, often making using only one fixed RM to train LLMs suboptimal. However, optimizing LLMs with multiple RMs simultaneously can incur a prohibitively high computational cost and lead to conflicting signals from different RMs that may degrade performance. To address these challenges, we introduce LASeR (Learning to Adaptively Select Rewards), which frames reward model selection as a multi-armed bandit problem, efficiently and iteratively training LLMs using multiple RMs by selecting the most well-suited RM for each instance. On commonsense and math reasoning tasks, we show that LASeR boosts iterative LLM training, improving the absolute average accuracy of Llama-3-8B over three datasets by $2.67\%$ over an ensemble of RM scores while also showing superior efficiency (e.g., a 2×2$\times$2 × speedup). Moreover, on WildChat (open-ended instruction-following tasks), LASeR leads to a $72.69\%$ AlpacaEval win rate over the RM score ensemble baseline. Extending to long-context generation, LASeR improves by $2.96$ F1 points (avg.) on single-document QA tasks and $2.97$ F1 points on few-shot learning over the RM score ensemble baseline with best-of-$n$ sampling. [^1] Code: https://github.com/duykhuongnguyen/LASeR-MAB
11footnotetext: Equal Contribution
@@ 1 Introduction
When comparing two responses, human preferences often differ depending on factors like the underlying task, who the annotators are (Santurkar et al., 2023; Ahmadian et al., 2024), and how preferences are elicited (Bansal et al., 2024). Therefore, models of preference data are also likely to differ and might include noise as well as any biases contained in the preference data used to train them. This can pose a problem when using such models as “reward models” (RMs) to align large language models (LLMs) to human preferences using reinforcement learning with human feedback (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022). Recent work has focused on aligning LLMs through iterative training, using reward models as proxies for human judgment (Gulcehre et al., 2023), leveraging the LLM to act as an implicit RM or judge (Yuan et al., 2024b; Chen et al., 2024b), or using the gold answer to compute a reward (Pang et al., 2024). Under this paradigm, there are three stages to training LLMs: (i) generating multiple responses to a query from an LLM; (ii) scoring the responses with an RM to create preference data with better and worse responses; and (iii) using model-generated preference data to further train the LLM. Note that for most domains, the gold reward is not readily available, making the quality of the RM or the degree to which it reflects human preferences (i.e., the gold reward) crucial to improving LLM performance. Indeed, several prior efforts aim to train new RMs that better reflect human preferences (Lambert et al., 2024).
@@
However, selecting one reward model to guide LLM training can be suboptimal for three main reasons: (1) A single RM may not be generalized to heterogeneous sets of examples. RMs are typically designed to reflect specific objectives and may be trained on offline preference datasets. Thus, an RM that performs well on one dataset or domain may not generalize effectively to others, leading to misaligned outputs across different tasks or domains (Kirk et al., 2023; Chen et al., 2024a; Casper et al., 2023; Gao et al., 2023). For instance, creativity plays a key role in evaluating the quality of a story, whereas correctness is more important in scoring math solutions. (2) RM performance leaderboards (e.g., Lambert et al. (2024)) that rely on human-annotated preferences can have unreliable rankings due to the presence of incorrect and ambiguous preferences (Yu et al., 2024; Hejna et al., 2023). (3) Lastly, over-optimization on one RM can lead to reward hacking problem (Skalse et al., 2022; Rafailov et al., 2024a), even resulting in downstream performance drops.
@@
To mitigate these issues, a prevalent approach is to ensemble multiple reward models (Coste et al., 2023; Eisenstein et al., 2023; Zhang et al., 2024; Ramé et al., 2024). However, these methods also come with significant challenges: as RMs are typically based on LLMs, training with multiple RMs often requires loading and managing several large models simultaneously, which can be computationally expensive, becoming infeasible as models increase in size. Moreover, aggregating multiple RM scores together is susceptible to noisy rewards or conflicting preferences from RMs, especially RMs that are not well-suited for the specific task (Rita et al., 2024). This, in turn, can degrade the quality of the preference data, leading to low-quality updates during training (Wang et al., 2024a). Finally, manually selecting a subset of RMs to combine is a labor-intensive process that involves training many different variants on a combinatorially large set of RM groupings. This underscores the need for more efficient and robust LLM optimization with multiple RMs.
@@
To fill this gap, we introduce Learning to Adaptively Select Rewards (LASeR), that, given a set of RMs, adaptively and efficiently selects a suitable RM for each instance by casting selection as a multi-armed bandit problem (Vermorel and Mohri, 2005; Audibert et al., 2009). Specifically, during training, the RM (arm) is chosen dynamically based on contextual information about the model’s performance and past interactions. The LLM is then fine-tuned based on the RM-annotated data, and the bandit’s parameters are updated accordingly to reflect the performance of the LLM after training on preference data annotated using selected RM (see Fig. 1). By design, LASeR’s adaptive instance-level RM selection (c.f. Sec. 3) addresses the three shortcomings of choosing one reward model: lack of generalization, unreliable rankings, and over-optimization. LASeR eliminates the need for manual tuning or oracle RM selection, with users only tasked with selecting a pool of RMs from a leaderboard like RewardBench (Lambert et al., 2024) without needing to know in advance which RM is best suited for a specific task. Moreover, unlike previous multi-RM methods that do not address conflicting signals from multiple RMs and require simultaneously loading and running multiple RMs (Ramé et al., 2024; Coste et al., 2023) or training an RM for a specific task (Quan, 2024), LASeR selects one RM at a time (Sec. 4), avoiding conflicts and improving efficiency.
@@
Empirically, we demonstrate the effectiveness of LASeR for iteratively training LLMs using multiple RMs on three broad domains: reasoning, instruction-following in text generation, and long-context understanding (Sec. 4.2). We show that on reasoning benchmarks such as StrategyQA (Geva et al., 2021) (testing commonsense reasoning) and GSM8K (Cobbe et al., 2021) (testing math reasoning), LASeR with Llama-3-8B improves absolute accuracy (averaged across 3 datasets) by $1.45\%$ over the best single RM and $2.67\%$ over an ensemble of RM scores. LASeR is also effective on general instruction-following: we show that using LASeR with four strong 7B RMs from RewardBench to finetune Llama-3-8B on a subset of WildChat (Zhao et al., 2024) beats LLMs trained with the best RM and with a baseline that sequentially selects RMs, with $56.34\%$ and $71.45\%$ win rates (respectively) on length-controlled AlpacaEval (Dubois et al., 2024). LASeR also beats the ensemble of RM scores baseline with $72.69\%$ win rates using Llama-3-8B. Moreover, our results show the effectiveness of LASeR’s RM selection strategy at inference-time on long-form generation tasks; on LongBench (Bai et al., 2022), we find LASeR beats the ensemble of RM scores baseline by $2.96$ F1 points on single-document QA tasks and $2.97$ F1 points on few-shot learning when using best-of-$n$ sampling for Llama-3-8B. Our analysis reveals that LASeR is more efficient than sequential multi-RM and RM score ensemble baselines in terms of training time (wall-clock hours) by a factor of 3×3$\times$3 ×, and 2×2$\times$2 ×, respectively, while being more robust to conflicting preferences (Sec. 5).
@@ 2 Related Work
Multiple Reward Ensembles. Recent work explores training LLMs using multiple rewards, often via ensembles (Ramé et al., 2024; Wu et al., 2024; Coste et al., 2023; Zhang et al., 2024; Wang et al., 2024b; Jang et al., 2023; Eisenstein et al., 2023). These methods typically aggregate or align scores across RMs, but suffer from inefficiency (both in terms of training and using multiple RMs) and conflicting rewards (Rita et al., 2024). In contrast, LASeR selects one pretrained RM per training step, avoiding these issues while outperforming ensemble-based baselines (Coste et al., 2023; Wang et al., 2024d). Other approaches train an RM from interpretable objectives using a human-annotated dataset (Wang et al., 2024c) or jointly train multiple task-specific RMs with a sparse router (Quan, 2024). Unlike these methods, LASeR instead uses off-the-shelf RMs to train on LLM-generated outputs, which has shown better performance (Ivison et al., 2024), and dynamically selects RMs from a leaderboard using a multi-armed bandit without any RM training or annotated datasets.
Iterative LLM Training. Standard RLHF pipelines rely on static human preference data (Ouyang et al., 2022; Bai et al., 2022; Touvron et al., 2023), which limits scalability by the size and quality of annotated preference data and the effectiveness of off-policy optimization (Xu et al., 2023; Xiong et al., 2024; Yuan et al., 2024b; Guo et al., 2024). Recent iterative methods improve this by generating feedback from model outputs, using gold labels (Singh et al., 2023; Pang et al., 2024), single RMs (Gulcehre et al., 2023), or self-judgment (Yuan et al., 2024b; Chen et al., 2024b). LASeR extends this by leveraging multiple RMs, avoiding pitfalls of single RM reliance (unreliable rankings, lack of generalization, and over-optimization), and reducing user burden in RM selection for a specific task (Huang et al., 2023).
Multi-Armed Bandits (MABs). MABs have been widely applied in ML for optimization and selection tasks (Chen et al., 2013; Li et al., 2010, 2018; Graves et al., 2017), including language models (Pasunuru et al., 2020; Krishnamurthy et al., 2024; Dwaracherla et al., 2024). LASeR uses MABs to dynamically select RMs during training, unlike prior work, which uses MABs to select the annotated samples for a single fixed RM (Dwaracherla et al., 2024) or uses MABs for model selection at test-time without RMs (Nguyen et al., 2024; Li, 2025). This allows LASeR to improve LLMs iteratively without gold labels or explicit optimization for evaluation metrics.
@@ 3 LASeR: Learning to Adaptively Select Rewards
First, we expand on the single-iteration training pipeline with a general reward function (Sec. 3.1). Then, in Sec. 3.2, we describe how LASeR dynamically selects an RM using MAB algorithms, assigning the RM for a given instance or batch. Finally, in Sec. 3.3, we describe cross-iteration training and how we update the MAB parameters. An illustration of LASeR is shown in Fig. 1.
Figure 1: Overview of LASeR. Given the query, the multi-armed bandit selects an RM depending on the underlying query and the bandit’s parameters (based on the usage of each RM and the expected MAB reward). At iteration $m$, the LLM generates multiple responses that are scored based on the selected RM for that query. These responses are ranked into preference pairs, which are then used to fine-tune the model. The same train loss $\mathcal{L}^{m}$ is used to update the parameters of the LLM as well as the MAB for the next iteration, making the entire pipeline iterative.
@@ 3.1 Training LLMs Using a Reward Function
Notation.   Following Yuan et al. (2024b) and Pang et al. (2024), we adopt an iterative training pipeline to finetune the LLM for $M$ iterations. Let $\pi_{m}$ be the LLM at iteration $m$; we assume that we start from an initial pretrained model $\pi_{0}$. Let $\mathcal{D}=\{x_{1},x_{2},\dots,x_{N}\}$ represent the training inputs, where $x_{i}$ is an input query or prompt. Corresponding to each input query $x_{i}$, we sample a set of $n$ responses from the LLM at the current $m^{\text{th}}$ iteration as $\mathbf{y}_{i}=\{y^{1}_{i},y^{2}_{i},\dots,y^{n}_{i}\}\sim\pi_{m}(y|x_{i})$. Let $R^{\star}\!:\!(y_{i}^{j}|x_{i})\rightarrow\mathbb{R}$ be a reward function that can score an LLM-generated response $y_{i}^{j}$ to a query $x_{i}$ based on how well it aligns with specific task objectives or instructions. Note that $R^{\star}(.)$ can be any reward function and may correspond to a single RM, one of the multiple RMs selected by the MAB (as in our case), or even the true reward.
@@
Generating Preference Pairs.   We evaluate each response $y_{i}^{j}$ using the reward function $R^{\star}(y_{i}^{j}|x_{i})$. By comparing the rewards assigned to different responses, we can form $P$ preference pairs $(y_{i}^{w},y_{i}^{l})$, where $y_{i}^{w}$ is preferred over $y_{i}^{l}$ if $R^{\star}(y_{i}^{w}|x_{i})>R^{\star}(y_{i}^{l}|x_{i})$, thereby building a preference dataset: [^2] Following Pang et al. (2024), we randomly sample $P\!=\!10$ pairs corresponding to each prompt $x_{i}$. For brevity, we omit this in the notation of $\mathcal{D}_{\mathrm{pref}}$; but in our setting $|\mathcal{D}_{\mathrm{pref}}|=P\times|\mathcal{D}|$.
$\mathcal{D}_{\mathrm{pref}}=\{(x_{i},y_{i}^{w},y_{i}^{l})\mid x_{i}\in\mathcal{D},R^{\star}(y_{i}^{w})>R^{\star}(y_{i}^{l}).$
@@
Training Loss Function ($\mathcal{L}^{m}$).   In each iteration, we fine-tune the model using the generated preference dataset $\mathcal{D}_{\mathrm{pref}}$, resulting in $M$ models $\pi_{1},\pi_{2},\dots,\pi_{M}$. Specifically, we update the model using the DPO loss (Rafailov et al., 2024b) for learning from the preference pairs, which is consistent with other work on iterative LLM training (Yuan et al., 2024b; Pang et al., 2024). In this work, we use the following loss functions:
@@
$\mathcal{L}^{m}_{\textsc{dpo}}(\pi_{m})\!=\!-\mathbb{E}_{\mathcal{D}_{\mathrm{pref}}}\left[\log\sigma\left(\beta\big{(}r_{i}^{w}-r_{i}^{l}\big{)}\right)\right];\mathcal{L}^{m}_{\textsc{nll}}(\pi_{m})\!=\!-\mathbb{E}_{\mathcal{D}_{\mathrm{pref}}}\left[\frac{\log\pi_{m}(y_{i}^{w}\mid x_{i})}{|y_{i}^{w}|}\right],$
@@
where $r_{i}^{w}=\log\frac{\pi_{m}(y_{i}^{w}\mid x_{i})}{\pi_{m-1}(y_{i}^{w}\mid x_{i})}$, $r_{i}^{l}=\log\frac{\pi_{m}(y_{i}^{l}\mid x_{i})}{\pi_{m-1}(y_{i}^{l}\mid x_{i})}$, $\pi_{m}$ and $\pi_{m-1}$ denotes the LLM in the current iteration $m$ and the previous iteration $m-1$ (used as the reference model in DPO loss). Following Yuan et al. (2024b), we use the standard DPO loss for instruction-finetuning. Following Pang et al. (2024), we use the NLL loss on the preferred responses as an additional regularizer for reasoning tasks, i.e., $\mathcal{L}^{m}\!=\!\mathcal{L}^{m}_{\textsc{dpo}}+\mathcal{L}^{m}_{\textsc{nll}}$. In Appendix C, we show that LASeR outperforms baselines irrespective of the choice of the loss function $\mathcal{L}^{m}$.